# Magnetic Flux Emergence in Binary Neutron Star Remnants

### v1

@@ I Introduction
During a binary neutron star (BNS) merger, the shearing motion between the stellar surfaces as they contact triggers turbulent instabilities, which help amplify the magnetic field to magnetar strengths within the first few milliseconds after merger [1, 2, 3, 4, 5, 6]. In the post-merger phase, this shear layer where the KHI is active settles to a region ${\sim}1~{}\mathrm{km}$ below the surface [7, 8], which is where the strongest fields are focused [5, 6]. Recent simulations suggest these strongly amplified fields can break out from the remnant star on relatively short timescales and help power short gamma-ray bursts (sGRBs) [9, 10]. However, it is not immediately clear how this breakout occurs; BNS remnants are typically stable against convection [7, 8, 11, 12], so this breakout cannot be purely hydrodynamical in nature.
A similar problem occurs in solar physics. Features such as coronal loops and prominences are direct consequences of strong magnetic fields emerging from the solar surface. These fields are generated by the solar dynamo, and they are propelled upward through the convective zone. Their progress then stalls inside the photosphere, which is stable against convection. A rather successful model for achieving emergence relies on magnetic buoyancy: as magnetic flux accumulates near the base of the photosphere, the magnetized region eventually becomes buoyant and accelerates upward [13].
@@
A similar mechanism might be responsible for strong post-merger fields, which emerge from a BNS and create magnetically-driven outflows that could potentially power relativistic jets. A strong toroidal magnetic field develops inside the shear layer (see, e.g., Refs. [14, 15]), which could become buoyant as it settles into hydrostatic equilibrium with the surrounding fluid. From the Newtonian ideal magnetohydrodynamics (MHD) equations, the velocity $\mathbf{v}$ of the flux tube evolves as
@@
$\rho^{\ast}\frac{D\mathbf{v}}{Dt}=\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}-\Delta\rho\mathbf{g},$ (2)
@@
Note the presence of the tension force $T\propto B^{2}/r$ on the right-hand side, which arises because we consider a toroidal field. This differs slightly from buoyancy in solar physics, which typically concerns relatively small sections of magnetized flux tubes. In the solar regime, a parallel-plane approximation is appropriate (e.g., [16, 17, 18, 13, 19, 20]), so there is no tension force at all. In the case of a BNS remnant, however, we are interested in the large-scale circumstellar toroidal field, and tension cannot be neglected. Recalling that $g(r)=Gm(r)/r^{2}$ for a spherically symmetric star, this implies that the field is only buoyant if
@@
Note that this condition only determines buoyancy at a given moment and does not necessarily mean a toroidal field inside a BNS remnant will emerge; similar to classic hydrodynamic buoyancy, a stably stratified atmosphere can impede magnetic buoyancy, and finite-size effects (e.g., drag) as the scale height shrinks can also affect buoyancy [21, 16]. Nevertheless, this condition suggests that magnetic buoyancy may be a possible mechanism for embedded magnetic fields in BNS remnants to emerge.
@@ II.1 Initial Data
We construct our model based on the results of Ref. [11]. As shown in Figures 3, 7, and 8 of Ref. [11], the remnant settles into a quasi-steady state approximately $10~{}\mathrm{ms}$ after merger. We perform nonlinear fits as functions of density (see Appendix A) to the temperature and lepton fraction profiles in order to compute a 1D finite-temperature equilibrium slice of the HS(DD2) EOS (hereafter DD2)[22, 23], which we then use to generate a static neutron star via the Tolman-Oppenheimer-Volkoff equations [24, 25]. Using a central density $\rho=2.2\times 10^{-3}~{}\textrm{M}_{\odot}^{-2}\approx 1.36\times 10^{15}~{}\mathrm{g}/\mathrm{cm}^{3}$, the resulting star has a mass of $M\approx 2.43~{}\textrm{M}_{\odot}$ and a radius $R\approx 8.35~{}\textrm{M}_{\odot}=12.3~{}\mathrm{km}$. Because our table only covers a finite density range, we define the edge of the star as the point dropping below $\rho=10^{-5}~{}\textrm{M}_{\odot}^{-2}\approx 6.18\times 10^{13}~{}\mathrm{g}/\mathrm{cm}^{3}$.
Because we are interested in the dynamics of buoyant magnetic structures, we add a toroidal flux tube of radius $s_{0}$ satisfying
@@
A realistic magnetic field in a merger is not completely axisymmetric, which can trigger additional magnetic buoyant instabilities beyond the simple form of magnetic buoyancy already discussed. Among the most important of these effects is the so-called Parker instability (or undular instability) [26, 27]. Consider a longitudinal perturbation (i.e., parallel to the field lines) to the density or velocity to a state otherwise in equilibrium. The field lines will bend and form a series of peaks and troughs. The fluid will attempt to slide along the field lines, which causes the peaks to become lighter and more buoyant while the troughs become heavier and thus fall. The Parker instability can often grow in a stratified medium where other forms of magnetic buoyancy cannot, which makes it an attractive prospect for enhancing buoyancy in our strongly stratified remnant profile.
@@
along the length of the tube where $A$ is a specified amplitude and $k$ is an integer determining the harmonic; because we keep the pressure fixed, this indirectly adds a perturbation to the temperature inside the flux tube. These longitudinal perturbations add a small amount of hydrodynamic buoyancy to the system, causing the field lines to bend and possibly triggering the Parker instability.
After constructing the flux tube, we enforce mechanical equilibrium by setting the tube’s pressure to $P^{\ast}=P_{0}-P_{B}$, where $P_{0}$ is the ambient fluid pressure and $P_{B}$ is magnetic pressure ($P_{B}=b^{2}/8\pi$ in CGS units, where $b^{\mu}$ is the magnetic field in the fluid’s rest frame). We then reset the fluid density to enforce thermal equilibrium with the (perturbed) ambient medium to create a buoyant force. In a true post-merger scenario, the thermal diffusion time scale is very long compared to the dynamical time scale, so calculating the density difference assuming an adiabatic process rather than an isothermal one may produce a more realistic flux tube. However, one can show that enforcing thermal equilibrium results in a larger density difference than an isentropic system, independent of the EOS, because $(\partial\rho/\partial P)_{T}>(\partial\rho/\partial P)_{s}$ is equivalent to requiring the heat capacities to satisfy $C_{P}/C_{V}>1$. Therefore, these tests tend to overestimate how buoyant a flux tube is and provide something of a best-case scenario for emergence.
@@
Name $\sqrt{b^{2}_{\mathrm{max}}}$ (G) $A$ $s_{0}$ ($\textrm{M}_{\odot}$) $t_{\mathrm{em}}$ (ms) $v_{\mathrm{em}}$
H7e17_5 $7.540\times 10^{17}$ $0.05$ $0.25$ 0.216 0.038
H7e17_5_B $7.540\times 10^{17}$ $0.05$ $1.0$ 0.201 0.013
H7e17_0 $7.540\times 10^{17}$ 00 $0.25$ 0.224 0.034
H1e17_5 $1.257\times 10^{17}$ $0.05$ $0.25$ N/A N/A
H1e17_0 $1.257\times 10^{17}$ 00 $0.25$ N/A N/A
H1e16_5 $1.257\times 10^{16}$ $0.05$ $0.25$ N/A N/A
H6e15_5 $6.283\times 10^{15}$ $0.05$ $0.25$ N/A N/A
H0_5 00 $0.05$ $0.25$ N/A N/A
For most of our runs, we consider a flux tube of inner radius $s_{0}=0.25~{}\textrm{M}_{\odot}\approx 369~{}\mathrm{m}$ and an outer radius $R$ matching the position of the peak temperature ($r=7.114~{}\textrm{M}_{\odot}$), which we take as a proxy for the position of the shear layer. We additionally set the width parameter $\sigma=0.3$ and add a small twist with $q=1$ to increase rigidity of the flux tube. The perturbation wavenumber is set to $k=18$.
Across our runs we vary the magnetic field strength and the perturbation amplitude. Based on the relativistic generalization of Eq. 5 (Eq. 24), only very weak fields ($B\lesssim 4\times 10^{12}~{}\textrm{G}$) are completely incapable of buoyancy. Due to the finite width of the flux tube and discretization effects, the necessary field strength for buoyancy can deviate somewhat from this ideal condition, but it serves as a useful guideline nonetheless. We test fields at $7.540\times 10^{17}~{}\textrm{G}$, $1.257\times 10^{17}~{}\textrm{G}$, $1.257\times 10^{16}~{}\textrm{G}$, and $6.283\times 10^{15}~{}\textrm{G}$. For every field strength, we perform a run with a perturbation amplitude $A=0.05$. We perform an additional test with $A=0$ for some cases to help investigate the role that the Parker instability plays in the perturbed runs. As a control, we also perform a run with no magnetic field but with an $A=0.05$ density perturbation; this will help disentangle the effects of ordinary hydrodynamic buoyancy from those caused by the Parker instability. To help track the evolution of the tube for this case, a passive tracer scalar $s$ is added with $s=1$ inside the tube and $s=0$ everywhere else.
@@
Since the shear layer is ${\sim}3~{}\mathrm{km}$ wide in a realistic BNS merger [8], a flux tube of $s_{0}=0.25~{}\textrm{M}_{\odot}$ may be much smaller in size than the toroidal field structure which develops in a realistic merger simulation. Because we may approximate the field as a superposition of weakly twisted flux tubes, tests with small flux tubes should be qualitatively correct when determining if a particular flux tube can emerge. Nevertheless, we do include one additional setup with the H7e17_5 configuration with a larger flux tube as well. For this run, we set $s_{0}=1.0~{}\textrm{M}_{\odot}\approx 1.48~{}\mathrm{km}$ and reduce the twist coefficient to $q=0.25$; since the twist scales linearly with the tube radius $s$, this is necessary to ensure the magnetic field remains primarily toroidal. We label this run H7e17_5_B to distinguish it from the case with a small flux tube.
We note that our strongest fields are unrealistically large for a BNS post-merger; high-resolution BNS simulations starting with realistic initial field strengths do not typically produce fields with peak strengths much in excess of $10^{17}~{}\textrm{G}$, and average strengths are typically ${\sim}10^{16}~{}\textrm{G}$ [3, 4, 28, 5, 14, 15]. Though no full-scale merger simulation has sufficient resolution to capture all scales of magnetic turbulence, local simulations confirm these findings [29, 30].
@@ II.2 Numerical Methods
We evolve the star with the GRMHD code AthenaK [31, 32, 33], which uses a second-order finite-volume method to evolve the fluid and upwind constrained transport to ensure that $\nabla\cdot\mathbf{B}=0$. For these tests, we choose WENOZ as our reconstruction scheme [34] and enable a first-order flux correction (FOFC) method [35]. The FOFC scheme greatly improves the stability of the evolution near the surface of the neutron star, particularly when the magnetization is high. Because we assume that the remnant has achieved steady state, we fix the spacetime to the initial TOV solution.
@@
Our computational domain is a wedge in the equatorial plane covering $r\in[6~{}\textrm{M}_{\odot},24~{}\textrm{M}_{\odot}]$, $\theta\in[\pi/2-\pi/9,\pi/2+\pi/9]$, and $\phi\in[-\pi/9,\pi/9]$, which amounts to full two oscillations for the $k=18$ perturbation described in Sec. II.1. We use a base domain of $600\times 180\times 180$ cells with a single refinement region over $r\in[6~{}\textrm{M}_{\odot},12\textrm{M}_{\odot}]$ for an effective radial resolution of $\Delta r=0.015~{}\textrm{M}_{\odot}\approx 22~{}\mathrm{m}$ inside the star. The angular resolution is sufficient that $r\Delta\theta\approx r\sin{\theta}\Delta\phi\approx\Delta r$ over the region of interest from approximately $7~{}\textrm{M}_{\odot}$ to $9~{}\textrm{M}_{\odot}$.
The upper boundary of $r$ is a “diode” outflow boundary (the radial momentum must be nonnegative), and the $\theta$ and $\phi$ boundaries are periodic. For the inner boundary of $r$, we keep the rest-mass density $\rho$ and pressure $P$ fixed, but we apply a reflecting boundary to the magnetic field $B^{i}$ and the radial fluid momentum density $\alpha\sqrt{\gamma}\rho Wv^{r}$, where $\gamma$ is the determinant of the spatial metric, $\alpha$ is the lapse, $v^{r}$ is the radial component of the fluid three-velocity, and $W=(1-v^{2})^{-1/2}$ is the Lorentz factor. This condition ensures that the density flux across the boundary is approximately zero. The $Wv^{\phi}$ and $Wv^{\theta}$ velocity components are also reflected across the boundary with the parity inverted. Though these components are parallel to the boundary, we found that inverting the parity improves stability; otherwise, small numerical errors can accumulate in the velocity, and a simple reflection can amplify these errors in a way that gives the star an unphysical rotation, while flipping the sign of the velocity components damps them.
@@
This work represents the first use of 3D EOS tables in AthenaK, which are implemented following the same prescription as in GR-Athena++ [36]. To approximate their effects on the remnant, we modify the full DD2 EOS table with the contributions of trapped neutrinos following the prescription in Ref. [37]. Therefore, the evolved species fraction is now the (electron) lepton fraction, $Y_{l}=Y_{e}+(n_{\nu_{e}}-n_{\nu_{\overline{e}}})/n_{b}$, which is advected along with the fluid flow. As $Y_{e}$ becomes a function of $Y_{l}$, $n_{b}$, and the temperature $T$, this prescription allows us to approximate the effects of trapped, fast-equilibriating neutrinos without needing to evolve them directly.

### v2

@@ I Introduction
During a binary neutron star (BNS) merger, the shearing motion between the stellar surfaces as they contact triggers turbulent instabilities, which help amplify the magnetic field to magnetar strengths within the first few milliseconds after merger Price and Rosswog (2006); Kiuchi et al. (2014, 2015, 2018); Palenzuela et al. (2022); Aguilera-Miret et al. (2023). In the post-merger phase, this shear layer where the Kelvin-Helmholtz instability is active settles to a region ${\sim}1~\mathrm{km}$ below the surface Kastaun and Galeazzi (2015); Hanauske et al. (2017), which is where the strongest fields are focused Palenzuela et al. (2022); Aguilera-Miret et al. (2023). Recent simulations suggest these strongly amplified fields can break out from the remnant star on relatively short timescales and help power short gamma-ray bursts (sGRBs) Most and Quataert (2023); Combi and Siegel (2023). However, it is not immediately clear how this breakout occurs; BNS remnants are typically stable against convection Kastaun and Galeazzi (2015); Hanauske et al. (2017); Radice and Bernuzzi (2023); Gao et al. (2026), so this breakout cannot be purely hydrodynamical in nature.
A similar problem occurs in solar physics. Features such as coronal loops and prominences are direct consequences of strong magnetic fields emerging from the solar surface. These fields are generated by the solar dynamo, and they are propelled upward through the convective zone. Their progress then stalls inside the photosphere, which is stable against convection. A rather successful model for achieving emergence relies on magnetic buoyancy: as magnetic flux accumulates near the base of the photosphere, the magnetized region eventually becomes buoyant and accelerates upward Archontis et al. (2004).
@@
A similar mechanism might be responsible for strong post-merger fields, which emerge from a BNS and create magnetically-driven outflows that could potentially power relativistic jets. A strong toroidal magnetic field develops inside the shear layer (see, e.g., Refs. Aguilera-Miret et al. (2022); Kiuchi et al. (2024)), which could become buoyant as it settles into hydrostatic equilibrium with the surrounding fluid. From the Newtonian ideal magnetohydrodynamics (MHD) equations, the velocity $\mathbf{v}$ of the flux tube evolves as
@@
$\rho^{\ast}\frac{D\mathbf{v}}{Dt}=\frac{1}{4\pi}\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}-\Delta\rho\mathbf{g},$ (2)
@@
Note the presence of the tension force $T\propto B^{2}/r$ on the right-hand side, which arises because we consider a toroidal field. This differs slightly from buoyancy in solar physics, which typically concerns relatively small sections of magnetized flux tubes. In the solar regime, a parallel-plane approximation is appropriate (e.g., Matsumoto et al. (1993); Longcope et al. (1996); Magara and Longcope (2001); Archontis et al. (2004); Murray et al. (2006); Toriumi et al. (2011)), so there is no tension force at all. In the case of a BNS remnant, however, we are interested in the large-scale circumstellar toroidal field, and tension cannot be neglected. Recalling that $g(r)=Gm(r)/r^{2}$ for a spherically symmetric star, this implies that the field is only buoyant if
@@
Note that this condition only determines buoyancy at a given moment and does not necessarily mean a toroidal field inside a BNS remnant will emerge; similar to classic hydrodynamic buoyancy, a stably stratified atmosphere can impede magnetic buoyancy, and finite-size effects (e.g., drag) as the scale height shrinks can also affect buoyancy Spruit et al. (1987); Matsumoto et al. (1993). Nevertheless, this condition suggests that magnetic buoyancy may be a possible mechanism for embedded magnetic fields in BNS remnants to emerge.
@@
Though most quantities in this paper are expressed in physical units, a number of formulae in Sec. III.1 and App. B instead use geometrized units with $G=c=1$ to maintain consistency with relativistic literature. In these scenarios, the solar mass $M_{\odot}$ is chosen as the relevant dimensional scale. However, any results obtained with these formulae quoted in the text are first converted to physical units.
@@ II.1 Initial Data
We construct our model based on the results of Ref. Radice and Bernuzzi (2023). As shown in Figures 3, 7, and 8 of Ref. Radice and Bernuzzi (2023), the remnant settles into a quasi-steady state approximately $10~\mathrm{ms}$ after merger. We perform nonlinear fits as functions of density (see Appendix A) to the temperature and lepton fraction profiles in order to compute a 1D finite-temperature equilibrium slice of the HS(DD2) EOS (hereafter DD2)Typel et al. (2010); Hempel and Schaffner-Bielich (2010), which we then use to generate a static neutron star via the Tolman-Oppenheimer-Volkoff (TOV) equations Tolman (1939); Oppenheimer and Volkoff (1939). Using a central density $\rho=2.2\times 10^{-3}~G^{-3}\textrm{M}_{\odot}^{-2}c^{6}\approx 1.36\times 10^{15}~\mathrm{g}/\mathrm{cm}^{3}$, the resulting star has a mass of $M\approx 2.43~\textrm{M}_{\odot}$ and a radius $R\approx 8.35~G\textrm{M}_{\odot}c^{-2}\approx 12.3~\mathrm{km}$. Because our table only covers a finite density range, we define the edge of the star as the point dropping below $\rho=10^{-5}~G^{-3}\textrm{M}_{\odot}^{-2}c^{6}\approx 6.18\times 10^{12}~\mathrm{g}/\mathrm{cm}^{3}$.
Because we are interested in the dynamics of buoyant magnetic structures, we modify the initial data, which is initially in hydrostatic equilibrium, by superimposing a toroidal flux tube of radius $s_{0}$ satisfying
@@
A realistic magnetic field in a merger is not completely axisymmetric, which can trigger additional magnetic buoyant instabilities beyond the simple form of magnetic buoyancy already discussed. Among the most important of these effects is the so-called Parker instability (or undular instability) Parker (1966); Acheson (1979). Consider a longitudinal perturbation (i.e., parallel to the field lines) to the density or velocity to a state otherwise in equilibrium. The field lines will bend and form a series of peaks and troughs. The fluid will attempt to slide along the field lines, which causes the peaks to become lighter and more buoyant while the troughs become heavier and thus fall. The Parker instability can often grow in a stratified medium where other forms of magnetic buoyancy cannot, which makes it an attractive prospect for enhancing buoyancy in our strongly stratified remnant profile.
@@
along the length of the tube where $A$ is a specified amplitude and $k$ is an integer determining the harmonic; because we keep the pressure fixed, this indirectly adds a perturbation to the temperature inside the flux tube. These longitudinal perturbations add a small amount of hydrodynamic buoyancy to the system, causing the field lines to bend and possibly seeding the Parker instability.
After superimposing the magnetic flux tube and applying the fluid perturbation, we enforce mechanical equilibrium with the TOV solution by setting the pressure inside the tube to $P^{\ast}=P_{0}-P_{B}$, where $P_{0}$ is the ambient fluid pressure and $P_{B}$ is magnetic pressure ($P_{B}=b^{2}/8\pi$ in CGS units, where $b^{\mu}$ is the magnetic field in the fluid’s rest frame). The density, which has already been perturbed, is then adjusted again to keep the (perturbed) temperature fixed. The composition of the fluid is also held constant during this process. Therefore, the fluid is in pressure equilibrium with the TOV solution and thermal equilibrium with the perturbed state, but it deviates slightly from thermal equilibrium with the TOV (leading to hydrodynamic buoyancy), and it cannot generally achieve hydrostatic equilibrium due to the magnetic field. In a true post-merger scenario, the thermal diffusion time scale is very long compared to the dynamical time scale, so calculating the density difference assuming an adiabatic process rather than an isothermal one may produce a more realistic flux tube. However, one can show that enforcing thermal equilibrium results in a larger density difference than an isentropic system, independent of the EOS, because $(\partial\rho/\partial P)_{T}>(\partial\rho/\partial P)_{s}$ is equivalent to requiring the heat capacities to satisfy $C_{P}/C_{V}>1$. Therefore, these tests tend to overestimate how buoyant a flux tube is and provide something of a best-case scenario for emergence.
@@
Name $\sqrt{b^{2}_{\mathrm{max}}}$ (G) AA $s_{0}$ (km) k $t_{\mathrm{em}}$ (ms) $v_{\mathrm{em}}(c)$
H7e17_5 $7.540\times 10^{17}$ $0.05$ $0.369$ 18 0.216 0.038
H7e17_5_B $7.540\times 10^{17}$ $0.05$ $1.48$ 18 0.201 0.013
H7e17_5_k9 $7.540\times 10^{17}$ $0.05$ $0.369$ 9 0.194 0.048
H7e17_0 $7.540\times 10^{17}$ 0 $0.369$ N/A 0.224 0.034
H1e17_5 $1.257\times 10^{17}$ $0.05$ $0.369$ 18 N/A N/A
H1e17_5_k9 $1.257\times 10^{17}$ $0.05$ $0.369$ 9 N/A N/A
H1e17_0 $1.257\times 10^{17}$ 0 $0.369$ N/A N/A N/A
H1e16_5 $1.257\times 10^{16}$ $0.05$ $0.369$ 18 N/A N/A
H6e15_5 $6.283\times 10^{15}$ $0.05$ $0.369$ 18 N/A N/A
H0_5 0 $0.05$ $0.369$ 18 N/A N/A
For most of our runs, we consider a flux tube of inner radius $s_{0}=0.25~G\textrm{M}_{\odot}c^{-2}\approx 369~\mathrm{m}$ and an outer radius $R$ matching the position of the peak temperature ($r=7.114~\textrm{M}_{\odot}$), which we take as a proxy for the position of the shear layer. We additionally set the width parameter $\sigma=0.3$ and add a small twist with $q=1~G^{-1}\textrm{M}_{\odot}^{-1}c^{2}\approx 0.677~\mathrm{km}^{-1}$ to increase rigidity of the flux tube. The perturbation wavenumber is set to $k=18$.
Across our runs we vary the magnetic field strength and the perturbation amplitude. Based on the relativistic generalization of Eq. 5 (Eq. 24), only very weak fields ($B\lesssim 4\times 10^{12}~\textrm{G}$) are completely incapable of buoyancy. Due to the finite width of the flux tube and discretization effects, the necessary field strength for buoyancy can deviate somewhat from this ideal condition, but it serves as a useful guideline nonetheless. We test fields at $7.540\times 10^{17}~\textrm{G}$, $1.257\times 10^{17}~\textrm{G}$, $1.257\times 10^{16}~\textrm{G}$, and $6.283\times 10^{15}~\textrm{G}$. For every field strength, we perform a run with a perturbation amplitude $A=0.05$. We perform an additional test with $A=0$ for some cases to help investigate the role that the Parker instability plays in the perturbed runs. As a control, we also perform a run with no magnetic field but with an $A=0.05$ density perturbation; this will help disentangle the effects of ordinary hydrodynamic buoyancy from those caused by the Parker instability. To help track the evolution of the tube for this case, a passive tracer scalar $X$ is added with $X=1$ inside the tube and $X=0$ everywhere else.
@@
Since the shear layer is ${\sim}3~\mathrm{km}$ wide in a realistic BNS merger Hanauske et al. (2017), a flux tube of $s_{0}=0.25~G\textrm{M}_{\odot}/c^{2}$ may be much smaller in size than the toroidal field structure which develops in a realistic merger simulation. Because we may approximate the field as a superposition of weakly twisted flux tubes, tests with small flux tubes should be qualitatively correct when determining if a particular flux tube can emerge. Nevertheless, we do include one additional setup with the H7e17_5 configuration with a larger flux tube as well. For this run, we set $s_{0}=1.0~G\textrm{M}_{\odot}/c^{2}\approx 1.48~\mathrm{km}$ and reduce the twist parameter to $q=0.25~G^{-1}\textrm{M}_{\odot}^{-1}c^{2}\approx 0.169~\mathrm{km}^{-1}$; since the twist scales linearly with the tube radius $s$, this is necessary to ensure the magnetic field remains primarily toroidal. We label this run H7e17_5_B to distinguish it from the case with a small flux tube.
We also consider two additional runs similar to H7e17_5 and H7e17_1 but with $k=9$, which we label as H7e17_5_k9 and H1e17_5_k9, respectively, to better understand any possible role of the Parker instability, which is wavenumber-dependent. In a plane-parallel approximation in Newtonian MHD, the maximum growth rate occurs at $\lambda\sim 10H_{P}$, where $\lambda$ is the wavelength and $H_{P}$ is the pressure scale height Horiuchi et al. (1988). This corresponds to $k\approx 8$ for our configuration, so we choose $k=9$ as it is the smallest perturbation that fits a full wavelength inside our domain. We acknowledge that our system is neither plane-parallel nor Newtonian, but we expect that the leading-order corrections to the growth rate from GR and a toroidal geometry should be similar across all wavenumbers. Note also that $k\approx 8$ is only the critical wavenumber near the initial location of the flux tube. As the flux tube rises, the scale height will rapidly decrease, so the critical wavenumber will increase. The “correct” perturbation for maximum growth, then, depends on the timescale for the Parker instability during the linear phase. If this growth rate is much shorter than the emergence timescale, the linear phase will be dominated by the initial conditions. At the other extreme, where the timescale is much longer than the emergence timescale, the Parker instability will not play a significant role in the dynamics, so the perturbation will be irrelevant.
We note that our strongest fields are unrealistically large for a BNS post-merger; high-resolution BNS simulations starting with realistic initial field strengths do not typically produce fields with peak strengths much in excess of $10^{17}~\textrm{G}$, and average strengths are typically ${\sim}10^{16}~\textrm{G}$ Kiuchi et al. (2015, 2018); Mösta et al. (2020); Palenzuela et al. (2022); Aguilera-Miret et al. (2022); Kiuchi et al. (2024). Though no full-scale merger simulation has sufficient resolution to capture all scales of magnetic turbulence, local simulations confirm these findings Obergaulinger et al. (2010); Zrake and MacFadyen (2013).
Note also that the remnant in Ref. Radice and Bernuzzi (2023) is differentially rotating, but we have neglected these effects by using a TOV solution. Though rotation can modify buoyancy and convective instabilities, it tends to inhibit these mechanisms Acheson and Gibbons (1978); Spruit (1999); Gao et al. (2026). In conjunction with our assumption of a flux tube in thermal (rather than isentropic) equilibrium, our results will tend toward overestimating the buoyancy of the system.
@@ II.2 Numerical Methods
We evolve the star with the GRMHD code AthenaK Stone et al. (2024); Zhu et al. (2025); Fields et al. (2025), which uses a second-order finite-volume method to evolve the fluid and upwind constrained transport to ensure that $\nabla\cdot\mathbf{B}=0$. For these tests, we choose WENOZ as our reconstruction scheme Borges et al. (2008) and enable a first-order flux correction (FOFC) method Lemaster and Stone (2009). The FOFC scheme greatly improves the stability of the evolution near the surface of the neutron star, particularly when the magnetization is high. Because we assume that the remnant has achieved steady state, we fix the spacetime to the initial TOV solution. Because the energy density inside the flux tube is reduced, there will be a gravitational mass deficit that should affect the metric. However, even for the H7e17_5_B case, this deficit is $\lesssim 2.2\%$ (and $\lesssim 0.2\%$ for H7e17_5), so this effect is a small perturbation that can safely be ignored.
@@
Our computational domain is a wedge in the equatorial plane covering $r\in[6~G\textrm{M}_{\odot}c^{-2},24~G\textrm{M}_{\odot}c^{-2}]\approx[8.86~\mathrm{km},35.44~\mathrm{km}]$, $\theta\in[\pi/2-\pi/9,\pi/2+\pi/9]$, and $\phi\in[-\pi/9,\pi/9]$, which amounts to full two oscillations for the $k=18$ perturbation described in Sec. II.1. We use a base domain of $600\times 180\times 180$ cells with a single refinement region over $r\in[6~G\textrm{M}_{\odot}c^{-2},12~G\textrm{M}_{\odot}c^{-2}]\approx[8.86~\mathrm{km},17.72~\mathrm{km}]$ for an effective radial resolution of $\Delta r=0.015~G\textrm{M}_{\odot}c^{-2}\approx 22~\mathrm{m}$ inside the star. The angular resolution is sufficient that $r\Delta\theta\approx r\sin{\theta}\Delta\phi\approx\Delta r$ over the region of interest from approximately $7~G\textrm{M}_{\odot}c^{-2}\approx 10~\mathrm{km}$ to $9~G\textrm{M}_{\odot}c^{-2}\approx 13~\mathrm{km}$.
The upper boundary of $r$ is a “diode” outflow boundary (the radial momentum must be nonnegative), which limits spurious fluxes generated at the level of the artificial atmosphere from inflowing back into the computational domain. The inner boundary of $r$ is designed to approximate a “stress-free” boundary. We first fix the rest-mass density $\rho$ and pressure $P$ to the initial TOV solution. The radial velocity $Wv^{r}$, where $W=(1-v^{2})^{-1/2}$ is the Lorentz factor and $v^{i}$ is the three-velocity, is then determined by setting $F_{\mathrm{in}}=-F_{\mathrm{out}}$ for the mass flux $F=\alpha\sqrt{\gamma}\rho Wv^{r}$, where $\gamma$ is the determinant of the spatial metric and $\alpha$ is the lapse, and solving for $(Wv^{r})_{\mathrm{in}}$. This behaves nearly like a typical reflection boundary, but it corrects for $\alpha\sqrt{\gamma}\rho$ being slightly larger on the inside of the star. The magnetic field and $Wv^{\phi}$ and $Wv^{\theta}$ are also reflected. Note, however, that we invert the parity for all components rather than only the radial component. In the event that the flux tube comes too close to the inner boundary, we found that a simple reflection that does not invert the parity of the tangential components of the magnetic field and velocity can cause small numerical errors to accumulate in the velocity and generate an unphysical rotation in the star. Through experimentation, we found that also flipping the tangential components damps these errors out and greatly improves stability. In practice the flux tube remains far from the radial boundary in all the runs presented here, so this condition has little effect on the evolution.
We exploit the spherical symmetry of the TOV solution and apply periodic boundaries in both the $\theta$ and $\phi$ directions. Though this means there are multiple copies of the flux tube along the polar direction, this is only an issue if the flux tube expands enough to interact with the polar boundary. So long as the flux tube is sufficiently resolved, this only happens some time after the flux tube emerges from the star; we end our runs before this occurs.
@@
This work represents the first use of 3D EOS tables in AthenaK, which are implemented following the same prescription as in GR-Athena++ Cook et al. (2025a). To approximate their effects on the remnant, we modify the full DD2 EOS table with the contributions of trapped neutrinos following the prescription in Ref. Perego et al. (2019). Therefore, the evolved species fraction is now the (electron) lepton fraction, $Y_{l}=Y_{e}+(n_{\nu_{e}}-n_{\nu_{\overline{e}}})/n_{b}$, which is advected along with the fluid flow. As $Y_{e}$ becomes a function of $Y_{l}$, $n_{b}$, and the temperature $T$, this prescription allows us to approximate the effects of trapped, fast-equilibriating neutrinos without needing to evolve them directly.